# The Neglected Tails of Vision-Language Models

### v2

@@ 4.2 REAL-Linear for Linear Classifier Training
REAL-Linear’s exceptional efficiency. REAL-Linear is significantly more efficient than the state-of-the-art REACT [24] and can be done using only academic-scale computing resources. Unlike REACT, which requires downloading the whole pretraining dataset and running VLMs to extract features for all pretraining images and texts, REAL-Linear processes only pretraining texts via string matching. As a result, it can process all LAION-400M captions in one hour, whereas REACT needs 250 GPU hours just for feature extraction. In addition, REACT’s contrastive training demands extensive resources for better performance, e.g., large batch sizes (4,096) and lengthy training (up to 256 GPU hours on 16 V100 GPUs). In contrast, our linear-probing approach trains in minutes on a modest (12GB) GPU. Table 3 compares the efficiency between REACT and REAL.
@@ 5 Experimental Results
We show that REAL establishes the new state-of-the-art for zero-shot recognition, outperforming existing prompting and retrieval-augmented methods across standard benchmarks. We ablate the design choices of REAL, revealing technical insights that contribute to its superior performance. We show that REAL can be combined with existing methods for even better results. Moreover, we show that REAL-Prompt improves image generation of rare concepts using text-to-image models like DALL-E 3 and SD-XL.
@@ 5.1 Experimental setup
| REAL-Prompt | 65.4 | 56.2 | 67.8 | 56.9 | |
| REAL-Linear | 67.8 | 58.9 | 72.5 | 56.0 | |
@@
| REAL-Prompt | 68.2 | 61.6 | 69.8 | 61.8 | |
| REAL-Linear | 69.8 | 64.8 | 76.2 | 63.6 | |
@@ Acknowledgements
We thank Tiffany Ling for helping polish figures.
@@ E More Ablations of REAL-Linear
| REAL-Linear (image-only) | 62.1 | 78.0 | 77.5 | 33.1 | 73.1 | 86.1 | 79.5 | 53.8 | 15.6 | 62.4 |
| REAL-Linear (cross-modal) | 65.9 | 78.8 | 84.1 | 29.6 | 74.0 | 89.5 | 81.4 | 61.5 | 51.5 | 67.8 |
@@
| REAL-Prompt | 63.6 | 76.6 | 82.7 | 18.0 | 64.0 | 88.8 | 81.0 | 59.9 | 57.5 | 65.8 |
| REAL-Linear (100) | 65.3 | 77.8 | 84.0 | 25.1 | 72.4 | 89.3 | 81.0 | 60.4 | 53.3 | 67.6 |
| REAL-Linear (500) | 65.9 | 78.8 | 84.4 | 29.6 | 74.0 | 89.5 | 81.4 | 61.5 | 51.5 | 68.5 |
@@ H REAL-Prompt for Generative Models
![[Uncaptioned image]](x7.png)
![[Uncaptioned image]](x8.png)
![[Uncaptioned image]](x9.png)
![[Uncaptioned image]](x10.png)
![[Uncaptioned image]](x11.png)
![[Uncaptioned image]](x12.png)
![[Uncaptioned image]](x13.png)
![[Uncaptioned image]](x14.png)
![[Uncaptioned image]](x15.png)
![[Uncaptioned image]](x16.png)
![[Uncaptioned image]](x17.png)
![[Uncaptioned image]](x18.png)
![[Uncaptioned image]](x19.png)
![[Uncaptioned image]](x20.png)
![[Uncaptioned image]](x21.png)
![[Uncaptioned image]](x22.png)

### v3

# The Neglected Tails in Vision-Language Models

@@ 1 Introduction
Popular multimodal systems like visual chatbots and generative models struggle with tailed concepts. An example is the night snake, a rare class found by our method on ImageNet.
@@ 4.2 REAL-Linear for Linear Classifier Training
REAL-Linear’s exceptional efficiency. REAL-Linear is significantly more efficient than the state-of-the-art REACT [24] and can be done using only academic-scale computing resources. Unlike REACT, which requires downloading the whole pretraining dataset and running VLMs to extract features for all pretraining images and texts, REAL-Linear processes only pretraining texts via string matching. As a result, it can process all LAION-400M captions in one hour, whereas REACT needs 250 GPU hours just for feature extraction. In addition, REACT’s contrastive training demands extensive resources to ensure performance, e.g., large batch size (4,096) and long training (256 GPU hours on 16 V100 GPUs). In contrast, our linear-probing approach trains in minutes on a modest GPU (12GB). Table 3 compares the efficiency between REACT and REAL.
@@ 5 Experimental Results
We show the state-of-the-art performance of REAL for zero-shot recognition, outperforming existing prompting and retrieval-augmented methods across standard benchmarks. We ablate the design choices of REAL, revealing technical insights that contribute to its superior performance. We show that REAL can be combined with existing methods for even better results. Moreover, we show that REAL-Prompt improves image generation of rare concepts using text-to-image models like DALL-E 3 and SD-XL.
@@ 5.1 Experimental setup
| REAL-Prompt | 65.4+0.6 | 56.2+1.0 | 67.8+2.1 | 56.9+4.4 | |
| REAL-Linear | 67.8+3.0 | 58.9+3.7 | 72.5+6.8 | 56.0+3.5 | |
@@
| REAL-Prompt | 68.2+0.2 | 61.6+0.6 | 69.8+1.2 | 61.8+3.4 | |
| REAL-Linear | 69.8+1.8 | 64.8+3.8 | 76.2+7.6 | 63.6+5.2 | |
@@ Acknowledgements
We thank Tiffany Ling for polishing figures. This work was partially supported by the University of Macau (SRG2023-00044-FST), and NSFC (No.62206256). Shubham Parashar acknowledges the support from the CSE Department at Texas A&M University.
@@ E More Ablations of REAL-Linear
| REAL-Linear (image-only) | 62.1-0.8 | 78.0+10.0 | 77.5-1.7 | 33.1+16.4 | 73.1+9.3 | 86.1-0.6 | 79.5-1.4 | 53.8-0.7 | 15.6-35.9 | 62.4-0.3 |
| REAL-Linear (cross-modal) | 65.9+3.0 | 78.8+10.8 | 84.1+7.9 | 29.6+12.9 | 74.0+10.2 | 89.5+2.8 | 81.4+0.5 | 61.5+6.2 | 51.5+0.0 | 67.8+6.1 |
@@
| REAL-Prompt | 63.6 | 76.6 | 82.7 | 18.0 | 64.0 | 88.8 | 81.0 | 59.9 | 57.5 | 65.8+3.1 |
| REAL-Linear (100) | 65.3 | 77.8 | 84.0 | 25.1 | 72.4 | 89.3 | 81.0 | 60.4 | 53.3 | 67.6+4.9 |
| REAL-Linear (500) | 65.9 | 78.8 | 84.4 | 29.6 | 74.0 | 89.5 | 81.4 | 61.5 | 51.5 | 68.5+5.8 |